OpenStudy (anonymous):

Given the equation 2 Square root of x minus 5 = 2, solve for x and identify if it is an extraneous solution. x = 6, solution is not extraneous x = 6, solution is extraneous x = 11, solution is not extraneous x = 11, solution is extraneous

OpenStudy (akashdeepdeb):

@meganemma1995 What exacrly do you mean by 2square root? :D

OpenStudy (anonymous):

2√x-5=2

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Okay so it is \[2 * \sqrt{x-5} = 2\] SQUARE BOTH SIDES 4(x-5) = 4 DIVIDE BOTH SIDES BY 4 x-5 = 1 ADD 5 TO BOTH SIDES x = 6 Understood? :)

OpenStudy (anonymous):

would it be extraneous or not? @AkashdeepDeb

OpenStudy (anonymous):

@AkashdeepDeb a extraneous is: a solution to an equation that does not fit the requirements of the original equation. and a non-extraneous equations is: an equation that does work

OpenStudy (anonymous):

was is extraneous?

OpenStudy (anonymous):

@xapproachesinfinity is this extraneous or not?

OpenStudy (xapproachesinfinity):

i said it is not did you read what i wrote i have given you the reason why

OpenStudy (anonymous):

this is a different equation though, let me see if I can figure it out and then tell me if im right or wrong?

OpenStudy (xapproachesinfinity):

Oh i didn't pay attention to the question i thought it was the one you asked before

OpenStudy (anonymous):

im going to say its not

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

and why it is not?

OpenStudy (anonymous):

because there was no 6 in the equation from the start? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

the simpliest why to know if it is or not is to check your answer in the equation does work or not

OpenStudy (xapproachesinfinity):

no that's not the reason

OpenStudy (anonymous):

I don't get it

OpenStudy (xapproachesinfinity):

do this to check \(\Large \tt\color{blueviolet}{2\sqrt{6-5}=2}\) we have to ask our selves is this true if it is true than the x=6 is not extraneous solution

OpenStudy (xapproachesinfinity):

it is a matter of checking if your answer satisfies the equation, yes?

OpenStudy (xapproachesinfinity):

So what do you think, actually phi has explained it in a better and sufficient way by now you should know how to do it^_^

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{2\sqrt{6-5}=2\sqrt{1}=2}\) so the solution works therefore it is not extraneous

OpenStudy (xapproachesinfinity):

you get it now?

OpenStudy (anonymous):

in order to figure out whether something is extraneous or not extraneous, is to basically plug the variable back into the equation after you have already solved for the variable, right?